DAQ Testing/20190115

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> DAQ Documentation Portal >> DAQ Testing >> DAQ Commissioning Notes

Previous Day of Testing << >> Next Day of Testing

January 15th, 2019 Testers: Cameron Clarke, Tao Ye, Bob

Goals

  • Get scalers working in HRSs
  • Use scalers for synchronization tests
  • Work on injector sync with fiber system and TS

Fiber Optics

  • The 4 fibers for the CH TS crate are verified to reach to the injector, but we need to run additional cables from the 2nd floor of CH to the first floor DAQ to allow for more signals.
  • We want to send some random DAQ V2F signals to injector (matching HRS crates scheme) and Brad Cumbia and Ed Jastrzembski's 8 channel sync scheme
  • Cables can be obtained to do this, but it will take some time to finalize it all

SIS3801 on HRSs

  • Follow up to run 4822 - the scaler was not being loaded in the HRS crate boot scripts, so no data was being read out
  • Loading the scaler fixes this and allows the scaler server Bob wrote to work (scread program reads the server stored data)
  • The LHRS had some issues where the baseline of the scaler values jumped around randomly and ultimately this was fixed by properly inserting the ROC into the VME backplane
  • Both HRSs have NIM scalers (CH has ECL scaler) in SIS3801 mode that we will use for now just for sync tests

Sync Tests

Tao and I took some runs with the DAQ trying to test the sync test on the 14th, and got some progress. The RHRS was not synched (due to sending Tstable instead of Tsettle to the RHRS)

  • Updates to DAQ:
    • We swapped the LNE for CH scaler from GMN2 VQWK gate made by HAPTB to Tsettle
    • And we sent Tsettle to both HRSs to be the trigger for all modules in each DAQ there
    • We generated a random DAC output from DAC12 in the HAPTB
      • Route this through a V2F module to get random frequency pulses instead of voltage
      • Make a Tsettle based gate that extends about half way through the helicity pulse length (240 Hz in Octet mode at this time due to injector studies)
      • And and not-And the gate with the V2F signal to make f1 and f2 to send to both LHRS and RHRS (had to find channels in the patch panel that were free, took some effort, but everything still works)
      • Plug the not-And into channel 31 and And into 32 in the HRS scalers, and put them into 4 and 5 respectively in the CH scaler - also put fr (frequency_reference) into the CH scaler channel 3 so that we can have the reference frequency there
      • Reproduce Bryan Moffit's Thesis Appendix C plots using Tao's ~apar/taoye/sync_monitor/Plot.C script
  • Runs:
    • 1017 and 1019 - tries this once, forgot to do random DAC right, and forgot to use Tsettle in CH crate
    • 1020 - got it right in LHRS and CH (scaler1 and scaler0), but now Tstable to RHRS causes problems (scaler2)
    • 1023 - used not-Tstable (Tsettle, but in a FIFO, not logic generator) - mostly fixes it in RHRS, but small delay
    • 1024 - used actual Tsettle and it all works now
    • 1027 and 1028 - double checking, both fail, probably due to restarting the ROC in CH to edit HAPTB timing to match 240 Hz (hopefully correct)
    • 1029 - also now change the HAPTB integrate time back - works again
    • 1030 - set it back to good - no errors of any form (pair sync error messages were appearing, probably due to Vqwk read/write errors)
    • 1031 - try again with 240 Hz haptb (and thus vqwk gate in GMN2 signal) - works! All Good!
  • We left the CH boot script VWQK and HAPTB settings in ~240 Hz mode
  • We set all LNEs in the HRSs to Tsettle (which will not work for proper asymmetry measurements in the VQWK gates - so this will necessitate more gate generators and patch panel signals)
  • We see (run 1031) that our scheme works just like in Bryan's thesis, and that it is possible to diagnose bad relative timing (run 1020 RHRS scaler - "scaler2")
  • For the plots corresponding to run 1020 and 1031 please see HAPLOG entry 3594, Bryan's Thesis see JLab' repository of theses

Next Steps

  • We will need to make sure that the LNE signals going into the ADCs and other modules in the HRSs are correct later, but for now the sync test works and everything is fine
  • We need to get the sync scheme working for the injector